GATE рдкреНрд░рд╢реНрди рдФрд░ рдЙрддреНрддрд░ рдХрд╛ рдЕрднреНрдпрд╛рд╕ рдХрд░реЗрдВ
рдкреНрд░: When water freezes its density 3961 05b5cc6a1e4d2b4197774cb08
5b5cc6a1e4d2b4197774cb08- 1decreasestrue
- 2remains constantfalse
- 3increasesfalse
- 4become zerofalse
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 1. "decreases"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: A) decreases Explanation: When water freezes its density decreases. Water at ordinary temperatures contracts and increases in density as it is cooled, like most substances. But at about 4┬░C it reaches a maximum density and then decreases in density as it approaches the freezing point.
рдкреНрд░: In a regular week, there are 5 working days and for each day, the working hours are 8. A man gets Rs. 2.40 per hour for regular work and Rs. 3.20 per hours for overtime. If he earns Rs. 432 in 4 weeks, then how many hours does he work for ? 3952 05b5cc74fe4d2b4197774fb35
5b5cc74fe4d2b4197774fb35- 1145false
- 2165false
- 3175true
- 4135false
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 3. "175"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: C) 175 Explanation: Given that, ┬аworking days = 5 ┬аworking hours = 8 ┬аA man get rupess per hour is Rs.2.40 ┬аSo in one day the man get total rupees is 2.40 x 8 = 19.2 ┬аSo in 5 days week the man get total rupees is 19.2 x 5 = 96 ┬аSo in 4 week the man get total rupees is 96 x 4 = 384 ┬аSo the man worked for = 160hours in 4 weeks ┬а ┬аBut given that the man earned Rs.432 ┬а ┬аHence remaning money is (432-384 = 48)which is earn by doing overtime work ┬а Overtime hours┬а = 48/3.20 = 15 So total worked hours is = 15 + 160 = 175.
рдкреНрд░: Re-arrange the following sentences into a meaningful paragraph. ┬а ( 3949 05b5cc6cde4d2b4197774e1f9
5b5cc6cde4d2b4197774e1f9- 1comes into effect from today. The bank's savings account rate was 3.5 per cent during March 2003 to May 2011 for all savings bank account customers (false
- 2The bank's decision comes two days ahead of RBI's monetary policy meet and is likely to trigger (false
- 3with deposits of above Rs 1 crore, it said in a statement. The two-tier savings bank account interest rate (false
- 4to 3.5 per cent, lowest in six years, a decision that will impact 90 per cent of its customers (E) The country's largest lender however will continue to pay 4 per cent interest on savings bank accounts (F) SBI today cut interest rate on savings bank account deposits up to Rs 1 crore by 0.5 per cent (G) rate war among peers. SBI's savings bank account base is of around Rs 9 lakh crore.true
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 4. "to 3.5 per cent, lowest in six years, a decision that will impact 90 per cent of its customers
(E) The country's largest lender however will continue to pay 4 per cent interest on savings bank accounts
(F) SBI today cut interest rate on savings bank account deposits up to Rs 1 crore by 0.5 per cent
(G) rate war among peers. SBI's savings bank account base is of around Rs 9 lakh crore."
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: D) FDBGECA Explanation:
рдкреНрд░: What is the next number in the given Number Series? 5, 24, 94, 279, ? 3942 05b5cc6b9e4d2b4197774d73e
5b5cc6b9e4d2b4197774d73e- 1587false
- 2554true
- 3489false
- 4499false
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 2. "554"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: B) 554 Explanation: The given number series follows a pattern that 5 x 5 тАУ 1 = 24 24 x 4 тАУ 2 = 94 94 x 3 тАУ 3 = 279 279 x 2 тАУ 4 = 554
рдкреНрд░: 8 couples (husband and wife) attend a dance show "Nach Baliye' in a popular TV channel ; A lucky draw in which 4 persons picked up for a prize is held, then the probability that there is atleast one couple will be selected is : 3941 05b5cc7d0e4d2b41977751218
5b5cc7d0e4d2b41977751218- 18/39false
- 215/39true
- 312/13false
- 4None of thesefalse
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 2. "15/39"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: B) 15/39 Explanation: P( selecting atleast one couple) = 1 - P(selecting none of the couples for the prize)┬а ┬а ┬а=┬а1-16C1├Ч┬а14C1├Ч12C1├Ч10C116C4=1539
рдкреНрд░: A man spend Rs. 810 in buying trouser at Rs. 70 each and shirt at Rs. 30 each. What will be the ratio of trouser and shirt when the maximum number of trouser is purchased ? 3931 05b5cc731e4d2b4197774f7b5
5b5cc731e4d2b4197774f7b5- 11:3false
- 22:1false
- 33:2true
- 42:3false
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 3. "3:2"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: C) 3:2 Explanation: Let us assume S as number of shirts and T as number of trousersGiven that each trouser cost = Rs.70 and that of shirt = Rs.30Therefore, 70 T + 30 S = 810=> 7T + 3S = 81......(1)T = ( 81 - 3S )/7 We need to find the least value of S which will make (81 - 3S) divisible by 7 to get maximum value of TSimplifying by taking 3 as common factor i.e, 3(27-S) / 7In the above equation least value of S as 6 so that 27- 3S becomes divisible by 7 Hence T = (81-3xS)/7 = (81-3x6)/7 = 63/7 = 9 Hence for S, put T in eq(1), we getS = 81-7(9)/3 = 81-63 / 3 = 18/3 = 6.The ratio of T:S = 9:6 = 3:2.
рдкреНрд░: In an election between two candidates K and L, the number of valid votes received by K exceeds those received by L by 15% of the total number of votes polled. If 20% of the votes polled were invalid and a total of 8720 votes were polled, then how many valid votes did L get ? 3928 05b5cc6d7e4d2b4197774e6fb
5b5cc6d7e4d2b4197774e6fb- 13241false
- 22457false
- 32642false
- 42834true
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 4. "2834"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: D) 2834 Explanation: Let the total number of votes polled in the election be 100k.Number of valid votes = 100k - 20% (100k) = 80kLet the number of votes polled in favour of K and L be a and b respectively.a - b = 15% (100k) => a = b + 15k=> a + b = b + 15k + bNow, 2b + 15k = 80k and hence b = 32.5kIt is given that 100k = 872032.5k = 32.5k/100k x 8720 = 2834 The number of valid votes polled in favour of L is 2834.
рдкреНрд░: In a queue, Amrita is 10th from the front while Mukul is 25th from behind and Mamta is just in the middle of the two. If there be 50 persons in the queue. What position does Mamta occupy from the front ? 3927 05b5cc6f8e4d2b4197774f068
5b5cc6f8e4d2b4197774f068- 114thfalse
- 216thfalse
- 318thtrue
- 420thfalse
- рдЙрддреНрддрд░ рджреЗрдЦреЗрдВрдЙрддреНрддрд░ рдЫрд┐рдкрд╛рдПрдВ
- Workspace
- SingleChoice
рдЙрддреНрддрд░ : 3. "18th"
рд╡реНрдпрд╛рдЦреНрдпрд╛ :
Answer: C) 18th Explanation: Number of persons between Amrita and Mukul = 50 - (10 + 25) = 15. Since Mamta lies in middle of these 15 persons, so Mamta`s position is 8th from Amrita i.e. 18th from the front.

